1.     Statistics Canada reported the following data for 2015:

Labor force: 20.4m, employment: 17m, working age Population: 26.7m. Calculate the number of people unemployed in 2015, unemployment rate and labor force participation rate. 


1
Expert's answer
2021-03-18T16:37:36-0400

i) Number of people unemployed =Labor Force "-" Employed population


"=20. 4M-17M"


"=3.4M"


ii) Unemployment Rate ;

"=\frac {Unemployed} {Labor Force} ×100"


"=\frac {3.4M}{20.4}=16.67\%"


iii) Labor Force Participation Rate ;is given by Labor Force divided by working-age population


"=\frac {Labor Force } {Working Age Population}"


="\frac{20.4}{26.7}×100 =76.4\%"